OpenStudy (anonymous):

Use partial Fractions to solve inegral 4y/ y^3-y^2+y-1

OpenStudy (kc_kennylau):

Hint: y-1 is one of the factors of y^3-y^2+y-1

OpenStudy (anonymous):

(y^2+1)(y-1) I keep trying to solve for A and B

OpenStudy (anonymous):

im doing it like Ay+B/9Y^2+1) + C/(y+1)

OpenStudy (kc_kennylau):

@satellite73 @Mertsj sorry my brain is stuck now idk y

OpenStudy (ranga):

Try: \[\Large \frac{ 4y }{ (y^2+1)(y-1)} = \frac{ Ay + B }{ y^2+1 } + \frac{ C }{ y-1 }\]

OpenStudy (anonymous):

that what I've been using to solve for A and B and C

OpenStudy (ranga):

you have a (y+1) under C. It should be (y-1).

OpenStudy (anonymous):

A+C = 0 B-A+4 C-B = 0

OpenStudy (ranga):

B-A = 4. Add all three and you will get C.

OpenStudy (anonymous):

A=-2 c=2 b=2